OpenStudy (anonymous):

Is x=5 a solution to -2x+10=0?

OpenStudy (anonymous):

LHS=-2(5)+10 =0 =RHS therefore, yes!

OpenStudy (anonymous):

coco right

OpenStudy (anonymous):

can I ask what LHS and RHS mean

OpenStudy (anonymous):

Rewrite as: [x/(x-5)] - 2 < 0 Rewrite with a common denominator: [x/(x-5)] - [2(x-5)/(x-5)] <0 Combine the numerators: to get: [x-2(x-5)]/(x-5) < 0 [-(x-10)]/[(x-5)] < 0 ---------------- The numerator is zero when x=10 The denominator is zero when x=5 Both of these are NOT part of the solution. -------------- Draw a number line; plot points x=5 and x=10 This breaks the number line into three intervals. --------------- Determine which interval(s) are part of the solution set: Pick a sample value from each interval and check it in the INEQUALITY, as follows: In the interval (-inf,5) pick x=0; then -(0-10)/(0-5)=-2 which is < 0 Therefore all x values in that interval are part of the solution set. --------------- In the interval (5,10) pick x = 8; then -(8-10)/(8-5)>0 Therefore no values in that interval are part of the solution set. ---------------- In the interval (10,inf) pick x=100; then -(100-10)/(100-5)<0 Therefore all x values in that interval are part of the solution set. ---------------------- Conclusion: The solution is all x-values in (-inf,5) OR (10,inf)

OpenStudy (anonymous):

LHS=left hand side RHS = right hand side

OpenStudy (anonymous):

Oh ok thank you